OpenStudy (adamaero):

Find the error. Calculus II's most confusing group work. No one in class got it today.

OpenStudy (adamaero):

I think the error is somewhere near: dy/dx = ky --> y = k(dx)

OpenStudy (adamaero):

any ideas SithsAndGiggles?

OpenStudy (anonymous):

\[\int\frac{dy}{y}=\ln|y|+C\not=\ln y+C\]

OpenStudy (adamaero):

So it was just missing abs val bars

OpenStudy (anonymous):

Essentially, yes. \[\begin{align*}\ln|y|&=kx+C\\ e^{\ln|y|}&=e^{kx+C}\\ |y|&=Ae^{kx} \end{align*}\] You can't have \(y=0\) because \(\ln 0\) is undefined, which means \(A\not=0\). Since \(|\text{anything}|\ge0\), then \(A>0\) exclusively.

OpenStudy (adamaero):

Cool beans